A and B can paint a certain house if A works 6 days and B works 12 days, or they can paint the house if A works 9 days and B works 8 days. How long would it take each working alone to paint the house?
Answer by richwmiller(9135)

12x+6y=1; 8x+9y=1
x = 1/20, y = 1/15
so x can do the job in 20 days
and y in 15 days if they worked alone.